OpenStudy (anonymous):

find the intervals on which the function is continuous. y= x+1 / (x^2 -15x +56)

OpenStudy (anonymous):

function is continuous at all points except where it is not defined. So, it is indirectly asking you to find the domain of the function. hint: any function gets undefined when the denominator gets zero.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

also, this is true only because the function is not a piecewise function.

OpenStudy (anonymous):

Yeah, I don't understand those. Like this one Find the limit of f(x) as x approaches 3 from the right f(x) { -2 if x < 3

OpenStudy (anonymous):

x+2 if 3 <= x <= 5 7 if x > 5

OpenStudy (anonymous):

from right means, coming towards "3", alittle more than 3, very close but not "3" so we use the second function to evaluate that

OpenStudy (anonymous):

How do I evaluate this ? x+2 if 3 <= x <= 5 The answer is 5 but I don't know how they got it.

OpenStudy (anonymous):

so, to begin with ou understood why we used the second definition?

OpenStudy (anonymous):

Because x approaches 3 from the right

OpenStudy (anonymous):

yes and so, x is greater than 3 now, it becomes a regular limit problem. you solve it by direct substitution

OpenStudy (anonymous):

\[ \lim_{x\to3^+}x+2=(3)+2=5 \]

OpenStudy (anonymous):

Ah, I see. I get it now! Thank you so much!

OpenStudy (anonymous):

yw